OpenStudy (anonymous):

if y=x+1 then whats the equation for y=f(ax) and y=af(x) if a=5?

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

My guess: f(x) = y = x + 1. f(ax) = f(5x) = (replace all 'x' with '5x') = 5x +1 af(x) = 5f(x) = (multiply f(x) by 5) = 5(x+1) = 5x + 5. Hope that's what you were looking for :P